OpenStudy (anonymous):

Simplify e^(3*ln(x))

OpenStudy (shubhamsrg):

let whole of it be y i.e. e^(3 ln x) = y take log to base e on both sides =>3 lnx = lny =>ln(x^3)=ln y or x^3=y = e^(3 lnx)

OpenStudy (shubhamsrg):

@mathslover nah,,you cant do that..that happens when powers add..

OpenStudy (anonymous):

Found the answer already, no need. X^3

mathslover (mathslover):

that i was going to say .. sorry

OpenStudy (shubhamsrg):

bah,,nevermind